OpenStudy (anonymous):

For 2x - 3y = -7, when x = 4, What is the value of y? give a medal to first that answers

OpenStudy (ajprincess):

Plug in the value of x in 2x-3y=-7. 2*4-3y=-7 8-3y=-7 can u find x nw @robertman228?

OpenStudy (anonymous):

is the answer -5

OpenStudy (ajprincess):

nope 8-3y=-7 8-3y+3y=-7+3y 8=-7+3y 8+7=-7+3y+7 15=3y 15/3=3y/3 y=5 Is it clear?

OpenStudy (anonymous):

alright got it, thanks

OpenStudy (ajprincess):

welcome:)